Kl(aq) + H2O(l) + KMnO4(aq) 166.0/mol 18.02g/mol 153.03g/mol l2(s) + MnO2(s) + KOH(aq) 253.80g/mol 86.97g/mol 56.11g/mol A.) Calculate the mass of manganese (lV) oxide that can be synthesized from 15.00 grams of potassium iodide. B.) Calculate the percent yield for this experiment if a mass of 1.982 grams of manganese (lV) oxide is produced. C.) Determine the percent error D.) compute for the excess amount of the excess reacta

1
Expert's answer
2021-10-29T02:30:26-0400

MnO2(s):KOH(aq)= 1:1

Molar mass of potassium iodide = 166.028

15/166.028= 0.09mol

Molar mass of Manganese IV Oxide = 86.94

= 86.94×0.09= 7.855g